# Impacts of axion cooling on the direct detection of supernova axions

### v1

@@ 1 Introduction
The axion is a hypothetical boson which is introduced to solve the strong CP problem in quantum chromodynamics [1, 2]. Since they are also a candidate for dark matter [3, 4, 5], massive efforts have been dedicated to experimental and astrophysical searches for axions [e.g. 6, 7, 8].
@@
Among astrophysical objects, core-collapse supernovae are one of the most luminous sources of axions because of the high temperature inside. Hence neutrinos and photons from SN 1987A, which appeared in the Large Magellanic Cloud, have been used to constrain the interaction between axions and the Standard Model particles. For example, axions work as an additional cooling process in the proto-neutron star (PNS), which can render the duration of the neutrino burst shorter than observed [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. In addition, if axions interact with photons, they can produce observable $\gamma$-rays during the propagation in the interstellar magnetic field [19, 20, 21, 22, 23]. The axion-photon conversion in the interstellar space may enable one to even detect the signature of the diffuse supernova axion background [24, 25, 26]. It is also notable that nearby massive stars before the core-collapse can produce a significant amount of axions that are converted into x-rays [27, 28, 29, 30].
@@
Apart from SN 1987A and distant stars, the Sun is another useful laboratory to study axions because of its proximity and stationarity. Axion helioscopes including CAST [31] and Tokyo Axion Helioscope [32] have observed the Sun to search for Solar axions through the axion-photon conversion in the strong magnetic field. Also, next-generation axion helioscopes such as IAXO [33] and TASTE [34] will lead to tighter constraints on the axion-photon interaction.
@@
Recently, the idea of direct detection of supernova axions with terrestrial detectors has been proposed [35, 36, 37]. In particular, ref. [35] proposed using the axion helioscopes. Since supernova axions are typically as energetic as $\sim 100$ MeV, they can be converted to $\gamma$-rays in strong magnetic fields. Therefore, if a $\gamma$-ray detector is installed on the other side of the x-ray detector for Solar axions, the axion helioscopes can be used as axion supernova-scopes. Since the axion burst from a supernova lasts only for $\sim 10$ s, the supernova-scopes should be aimed at the progenitor star before its explosion. Ref. [35] proposed to use the presupernova neutrino alert [38, 39, 40, 41] to target the supernova-scopes. Although the directional sensitivity of the inverse $\beta$-decay is weak, the unique identification of the progenitor is still possible [42, 43].
@@
In refs. [35, 37], it is estimated that $\mathcal{O}(1)$ axion events from $\alpha$ Orionis (Betelgeuse) can be detected by an improved version of IAXO, if the axion-photon coupling constant is as high as the upper limit obtained from SN 1987A neutrinos. However, when the axion luminosity is close to the neutrino luminosity, the proto-neutron star cooling is significantly accelerated [14, 15, 16, 17]. As a result, the axion luminosity could be overestimated if one adopts the post-process technique. In this study, we perform long-term supernova simulations coupled with the axion emission to consider the effect of the additional cooling on the supernova axion signal.
@@ 2 Method
We perform general-relativistic neutrino radiation hydrodynamics simulations in spherically symmetric geometry by employing the GR1D code [44, 45]. The code solves the neutrino transport with the truncated moment formalism [46, 47]. We use 18 logarithmically spaced energy bins from 1.0 to 280.5 MeV. We utilize the DD2 equation of state [48, 49], which is based on the relativistic mean field approach. The neutrino reaction rates are calculated with the NuLib routine [45]. As the initial condition, we adopt a $9.6M_{\odot}$ star with zero-metalicity provided by A. Heger (2016, private communication). The progenitor is chosen because the star successfully explodes even in the one-dimensional configuration without enhancing the neutrino heating rate by hand [50, 51]. The outer boundary for the simulations is located at the radius of $r_{\mathrm{out}}=5000$ km.
The original version of GR1D is optimized to simulate the accretion phase, which is $\lesssim 1$ s after the core bounce. Since the axion burst can continue after the accretion phase, we perform simulations until the post-bounce time $t_{\mathrm{pb}}=20$ s. In order to realize the long-term simulations, we modify the spatial resolution to resolve the steep density gradient at the surface of the PNS. The number of the grid points is $N_{r}=300$. In addition, we extend tables for thermodynamic quantities in the code because they can exceed the table in GR1D during the long-term simulations. See ref. [51] for details on the modification.
We calculate the axion emissivity during the simulations. We consider the Kim-Shifman-Vainshtein-Zakharov (KSVZ) axions [52, 53] as the benchmark model. Axions can interact with nucleons and photons as described by the Lagrangian [54]
@@
where $g_{aN}$ is the axion-nucleon coupling, $g_{a\gamma}$ is the axion-photon coupling, $N$ is nucleons, $a$ is axions, $F_{\mu\nu}$ is the electromagnetic tensor and $\tilde{F}_{\mu\nu}$ is its dual, and $m_{\mathrm{N}}$ is the nucleon mass. The coupling constants for KSVZ axions are given as $g_{aN}=C_{aN}m_{N}/f_{a}$ and $g_{a\gamma}=\alpha/\pi f_{a}$, where $m_{a}$ is the axion mass, $f_{a}$ is the axion decay constant, $C_{ap}=-0.47(3)$, and $C_{an}=-0.02(3)$ [55]. The axion mass and the decay constant have the relationship
@@
where $m_{u}$, $m_{d}$, $m_{\pi}$ are the mass of up quarks, down quarks, and neutral pions, respectively, and $f_{\pi}$ is the pion decay constant. We adopt the mass range of $m_{a}=3$–11 meV in our simulations, considering the SN 1987A upper limit $m_{a}\lesssim 5$–15 meV [15]. Table 1 is the list of our models. As shown in the table, we perform six simulations with axions and a simulation without axions.
@@
where $N$ and $N^{\prime}$ are nucleons. Within the one-pion exchange (OPE) approximation, the axion emissivity in vacuum is given as [56, 57, 15]
@@
where $f\approx 1$ is the pion-nucleon fine structure constant [58], $T$ is the temperature, $\bar{\xi}$ is a parameter to represent the degeneracy [57], $y_{n}$ and $y_{p}$ are the nucleon degeneracy parameters, and the function $I(y_{1},\,y_{2})$ is a fitting formula developed in ref. [56]. Taking into account the many-body effect that becomes important in the high density matter, we adopt in the simulations the axion emission rate [13, 14]
@@
The OPE treatment for the nucleon bremsstrahlung has been often adopted in literature to estimate the axion emission from supernovae [e.g. 13, 14, 17]. However, treatments beyond the OPE are recently discussed [e.g. 59, 15]. In ref. [15], the effects of the nonzero pion mass [60], the two-pions exchange [61], the effective nucleon mass, and the nucleon multiple scatterings [62, 63] are considered to calculate the axion emission from a supernova. They found that the axion luminosity is suppressed by an order of magnitude if these corrections are taken into account. In addition, it has been pointed out that supernova axions can be produced through other channels such as the pion-induced reaction ($\pi^{-}p\rightarrow na$) [57, 13, 64, 16] and a large number of reactions involved with strange matter [65]. These additional channels will enhance the axion emissivity. In particular, refs. [64, 16] discuss that the pion-induced reactions produce high-energy axions, which could be detected by future water-Cherenkov detectors through the axion-nucleon interaction. In order to predict the axion luminosity from supernovae accurately, it is desirable to perform simulations that take into account these corrections beyond the OPE and the additional axion production channels. However, since our motivation here is to investigate the importance of the axion cooling effect on the axion detectability, we leave this task for future work.
@@ 3.2 Expected event number of axions
An axion helioscope [66] is an instrument to detect Solar axions. A strong magnetic field is imposed in a cylindrical hollow to convert axions into x-rays through the axion-photon coupling. An x-ray detector is installed at the end of the cylinder to detect the signal. Axion helioscopes have been used to obtain tight upper limits on the axion-photon coupling $g_{a\gamma}$. For example, the CAST experiment reported an upper limit $g_{a\gamma}<0.66\times 10^{-10}$ GeV-1 at 95% confidence level [31].
As proposed in ref. [35], axion helioscopes can be used to detect supernova axions in principle. One of differences between Solar axions and supernova axions is their energy. Since supernova axions are as energetic as 10–100 MeV, they are converted into $\gamma$-rays by the magnetic field. In order to detect supernova axions with axion helioscopes, one should install an additional $\gamma$-ray detector. Another difference is that, whereas the Sun is a static object, supernova explosions are transient events. It is thus necessary that the axion helioscope is pointed to the supernova event before its neutrino burst is observed. This requires the identification of the progenitor star before the explosion, which could be possible through a presupernova neutrino alert [38, 39, 40, 42, 43, 41]. It is estimated that presupernova neutrinos can provide the alert $\mathcal{O}(1)$ hours before the core bounce for a massive star closer than 1 kpc, with detectors including the Jiangmen Underground Neutrino Observatory (JUNO) [67, 43], Super-Kamiokande [68, 69, 70], Hyper-Kamiokande [71], KamLAND [72, 73, 70], the Deep Underground Neutrino Experiment (DUNE) [74], and SNO+ [75].
The axion-photon conversion probability in an axion helioscope can estimated as [66, 76]
@@
Figure 5: The asymptotic event number of axions as a function of the axion luminosity $L_{a}$ at $t_{\mathrm{pb}}=1$ s. The left panel is based on the self-consistent simulations and the right panel is based on the post-processing. Each curve represents the results with different distances to the supernova event. The IAXO upgraded scenario is assumed as the detector. The vertical line at $L_{a}=3\times 10^{52}$ erg s-1 is an upper limit on the axion luminosity based on the SN 1987A neutrino burst [77].
@@
where $A$ is the cross section of the helioscope, $d$ is the distance to the supernova event, $\dot{N}_{a}=L_{a}/\langle\omega\rangle$ is the axion number produced in the supernova per unit time, and $\langle\omega\rangle\approx 2.3T$ is the axion mean energy. The cross section $A$, the magnetic field $B$, and the length $L$ depend on the detectors. In this study, we adopt values for the IAXO upgraded scenario shown in ref. [33], where $B\sim 35$ kG, $L=22$ m, and $A=3.9$ m2.
Figure 4 shows the cumulative event number of axions from Betelgeuse, which is located at $d\approx 168$ pc [78]. The solid curves are the results for the models with the axion cooling effect, and the broken curves are the results obtained by the post-processing of the reference model without the axion cooling. One can find that, when the axion cooling is not considered, the event number exceeds one if $m_{a}\gtrsim 9$ meV. On the other hand, when the axion cooling is taken into account, the cumulative event number saturates earlier and its asymptotic value is significantly reduced. As a result, the expected event number does not exceed one even if axions are as heavy as $m_{a}=11$ meV. The reduction in the event number stems from the reduction in the axion luminosity, which is attributed to the lower temperature in the supernova core shown in figure 3.
@@
HD 149757 $\zeta$ Ophiuchi O 112(2) [79] 20.0 [81]
HD 129056 $\alpha$ Lupi B 143(3) [79] $10.1\pm 1.0$ [82]
@@
Table 2: The list of nearby stars heavier than $10M_{\odot}$ and closer than 200 pc [42].
@@
The axion luminosity has an upper limit based on the neutrino burst from SN 1987A. If the axion luminosity dominates over the neutrino luminosity, the duration of the SN 1987A neutrino burst should be shorter than observed. This energy-loss argument has been applied to various exotic particle models to obtain constraints on the interaction between the new particle and a Standard Model particle. In many cases, the criterion of $L_{a}<3\times 10^{52}$ erg s-1 at $t_{\mathrm{pb}}=1$ s is adopted for this purpose [e.g. 77, 84, 85]. The vertical line in figure 5 indicates the SN 1987A limit. In our case, the models with $m_{a}\leq 9$ meV satisfies the condition, regardless of whether the axion cooling is considered or not. If we respect this upper limit on the axion luminosity, in the post-process framework, the axion signal is detectable if the star is closer than $d\sim 150$ pc. On the other hand, if we consider the axion cooling, the star should be closer than $d\sim 100$ pc for us to detect axions from it.

### v2

@@ 1 Introduction
The axion is a hypothetical boson which is introduced to solve the strong CP problem in quantum chromodynamics [1978PhRvL..40..223W, 1978PhRvL..40..279W]. Since they are also a candidate for dark matter [1983PhLB..120..133A, 1983PhLB..120..127P, 1983PhLB..120..137D], massive efforts have been dedicated to experimental and astrophysical searches for axions [e.g. 1978JETPL..27..502V, 2021ARNPS..71..225C, 2024arXiv240113728C, 2025PhyR.1117....1C].
@@
Among astrophysical objects, core-collapse supernovae are one of the most luminous sources of axions because of the high temperature inside. Hence neutrinos and photons from SN 1987A, which appeared in the Large Magellanic Cloud, have been used to constrain the interaction between axions and the Standard Model particles. For example, axions work as an additional cooling process in the proto-neutron star (PNS), which can render the duration of the neutrino burst shorter than observed [1988PhRvL..60.1797T, PhysRevD.39.1020, 1988PhRvL..60.1793R, 1990PhRvL..65..960E, 1997PhRvD..56.2419K, 2016PhRvD..94h5012F, 2019JCAP...10..016C, 2021PhRvD.104j3012F, PhysRevD.106.063019, 2023APh...15102855F, 2024PhRvD.109b3001L]. In addition, if axions interact with photons, they can produce observable $\gamma$-rays during the propagation in the interstellar magnetic field [1996PhRvL..77.2372G, 1996PhLB..383..439B, 2015JCAP...02..006P, 2017PhRvD..96e1701K, 2024PhRvL.133u1002M]. The axion-photon conversion in the interstellar space may enable one to even detect the signature of the diffuse supernova axion background [2011PhRvD..84j3008R, 2020PhRvD.102l3005C, 2022PhRvD.105f3028C]. It is also notable that nearby massive stars before the core-collapse can produce a significant amount of axions that are converted into x-rays [1995PhLB..344..245C, 2021PhRvL.126c1101X, 2022PhRvD.106l3019X, 2022PhRvD.105b3020M].
@@
Apart from SN 1987A and distant stars, the Sun is another useful laboratory to study axions because of its proximity and stationarity. Axion helioscopes including CAST [2017NatPh..13..584A] and Tokyo Axion Helioscope [2008PhLB..668...93I] have observed the Sun to search for Solar axions through the axion-photon conversion in the strong magnetic field. Also, next-generation axion helioscopes such as IAXO [2019JCAP...06..047A] and TASTE [2017JInst..12P1019A] will lead to tighter constraints on the axion-photon interaction.
@@
Recently, the idea of direct detection of supernova axions with terrestrial detectors has been proposed [2020JCAP...11..059G, 2022PhLB..82937137A, 2025arXiv250219476C]. In particular, ref. [2020JCAP...11..059G] proposed using the axion helioscopes. Since supernova axions are typically as energetic as $\sim 100$ MeV, they can be converted to $\gamma$-rays in strong magnetic fields. Therefore, if a $\gamma$-ray detector is installed on the other side of the x-ray detector for Solar axions, the axion helioscopes can be used as axion supernova-scopes. Since the axion burst from a supernova lasts only for $\sim 10$ s, the supernova-scopes should be aimed at the progenitor star before its explosion. Ref. [2020JCAP...11..059G] proposed to use the presupernova neutrino alert [2004APh....21..303O, 2017ApJ...851....6P, 2020ARNPS..70..121K, 2021NJPh...23c1201A] to target the supernova-scopes. Although the directional sensitivity of the inverse $\beta$-decay is weak, the unique identification of the progenitor is still possible [2020ApJ...899..153M, 2020JCAP...05..049L].
@@
In refs. [2020JCAP...11..059G, 2025arXiv250219476C], it is estimated that $\mathcal{O}(1)$ axion events from $\alpha$ Orionis (Betelgeuse) can be detected by an improved version of IAXO, if the axion-photon coupling constant is as high as the upper limit obtained from SN 1987A neutrinos. However, when the axion luminosity is close to the neutrino luminosity, the proto-neutron star cooling is significantly accelerated [2016PhRvD..94h5012F, 2019JCAP...10..016C, 2021PhRvD.104j3012F, PhysRevD.106.063019]. As a result, the axion luminosity could be overestimated if one adopts the post-process technique. In this study, we perform long-term supernova simulations coupled with the axion emission to consider the effect of the additional cooling on the supernova axion signal.
@@ 2 Method
We perform general-relativistic neutrino radiation hydrodynamics simulations in spherically symmetric geometry by employing the GR1D code [2010CQGra..27k4103O, 2015ApJS..219...24O]. The code solves the neutrino transport with the truncated moment formalism [2011PThPh.125.1255S, 2013PhRvD..87j3004C]. We use 18 logarithmically spaced energy bins from 1.0 to 280.5 MeV. We modify and optimize GR1D for long-term simulation [2021PTEP.2021b3E01M, 2025PASJ...77..127M]. We utilize the DD2 equation of state [2005PhRvC..71f4301T, 2010NuPhA.837..210H], which is based on the relativistic mean field approach. The neutrino reaction rates are calculated with the NuLib routine [2015ApJS..219...24O]. As the initial condition, we adopt a $9.6M_{\odot}$ star with zero-metalicity provided by A. Heger (2016, private communication). The progenitor is chosen because the star successfully explodes even in the one-dimensional configuration without enhancing the neutrino heating rate by hand [2015ApJ...801L..24M, 2021PTEP.2021b3E01M]. The outer boundary for the simulations is located at the radius of $r_{\mathrm{out}}=5000$ km.
The original version of GR1D is optimized to simulate the accretion phase, which is $\lesssim 1$ s after the core bounce. Since the axion burst can continue after the accretion phase, we perform simulations until the post-bounce time $t_{\mathrm{pb}}=20$ s. In order to realize the long-term simulations, we modify the spatial resolution to resolve the steep density gradient at the surface of the PNS. The number of the grid points is $N_{r}=300$. In addition, we extend tables for thermodynamic quantities in the code because they can exceed the table in GR1D during the long-term simulations. See ref. [2021PTEP.2021b3E01M] for details on the modification.
We calculate the axion emissivity during the simulations. We consider the Kim-Shifman-Vainshtein-Zakharov (KSVZ) axions [1979PhRvL..43..103K, 1980NuPhB.166..493S] as the benchmark model. Axions can interact with nucleons and photons as described by the Lagrangian [1989PhRvD..40..652C]
@@
where $g_{aN}$ is the axion-nucleon coupling, $g_{a\gamma}$ is the axion-photon coupling, $N$ is nucleons, $a$ is axions, $F_{\mu\nu}$ is the electromagnetic tensor and $\tilde{F}_{\mu\nu}$ is its dual, and $m_{\mathrm{N}}$ is the nucleon mass. The coupling constants for KSVZ axions are given as $g_{aN}=C_{aN}m_{N}/f_{a}$ and $g_{a\gamma}=\alpha/\pi f_{a}$, where $m_{a}$ is the axion mass, $f_{a}$ is the axion decay constant, $C_{ap}=-0.47(3)$, and $C_{an}=-0.02(3)$ [2016JHEP...01..034D]. The axion mass and the decay constant have the relationship
@@
where $m_{u}$, $m_{d}$, $m_{\pi}$ are the mass of up quarks, down quarks, and neutral pions, respectively, and $f_{\pi}$ is the pion decay constant. We adopt the mass range of $m_{a}=3$–11 meV in our simulations, considering the SN 1987A upper limit $m_{a}\lesssim 5$–15 meV [2019JCAP...10..016C]. Table 1 is the list of our models. As shown in the table, we perform six simulations with axions and a simulation without axions.
@@
where $N$ and $N^{\prime}$ are nucleons. Within the one-pion exchange (OPE) approximation, the axion emissivity in vacuum is given as [1988PhRvD..38.2338B, 1995PhRvD..52.1780R, 2019JCAP...10..016C]
@@
where $f\approx 1$ is the pion-nucleon fine structure constant [1998ApJ...507..339H], $T$ is the temperature, $\bar{\xi}$ is a parameter to represent the degeneracy [1995PhRvD..52.1780R], $y_{n}$ and $y_{p}$ are the nucleon degeneracy parameters, and the function $I(y_{1},\,y_{2})$ is a fitting formula developed in ref. [1988PhRvD..38.2338B]. Taking into account the many-body effect that becomes important in the high density matter, we adopt in the simulations the axion emission rate [1997PhRvD..56.2419K, 2016PhRvD..94h5012F]
@@
The OPE treatment for the nucleon bremsstrahlung has been often adopted in literature to estimate the axion emission from supernovae [e.g. 1997PhRvD..56.2419K, 2016PhRvD..94h5012F, PhysRevD.106.063019]. However, treatments beyond the OPE are recently discussed [e.g. 2018JHEP...09..051C, 2019JCAP...10..016C]. In ref. [2019JCAP...10..016C], the effects of the nonzero pion mass [2009NuPhA.828..439S], the two-pions exchange [1989PhLB..219..507E], the effective nucleon mass, and the nucleon multiple scatterings [1991PhRvL..67.2605R, 1996PhRvL..76.2621J] are considered to calculate the axion emission from a supernova. They found that the axion luminosity is suppressed by an order of magnitude if these corrections are taken into account. In addition, it has been pointed out that supernova axions can be produced through other channels such as the pion-induced reaction ($\pi^{-}p\rightarrow na$) [1995PhRvD..52.1780R, 1997PhRvD..56.2419K, 2021PhRvL.126g1102C, 2021PhRvD.104j3012F] and a large number of reactions involved with strange matter [2024PhRvL.133l1002C]. These additional channels will enhance the axion emissivity. In particular, refs. [2021PhRvL.126g1102C, 2021PhRvD.104j3012F] discuss that the pion-induced reactions produce high-energy axions, which could be detected by future water-Cherenkov detectors through the axion-nucleon interaction. In order to predict the axion luminosity from supernovae accurately, it is desirable to perform simulations that take into account these corrections beyond the OPE and the additional axion production channels. However, since our motivation here is to investigate the importance of the axion cooling effect on the axion detectability, we leave this task for future work.
@@ 3.2 Expected event number of axions
An axion helioscope [1983PhRvL..51.1415S] is an instrument to detect Solar axions. A strong magnetic field is imposed in a cylindrical hollow to convert axions into x-rays through the axion-photon coupling. An x-ray detector is installed at the end of the cylinder to detect the signal. Axion helioscopes have been used to obtain tight upper limits on the axion-photon coupling $g_{a\gamma}$. For example, the CAST experiment reported an upper limit $g_{a\gamma}<0.66\times 10^{-10}$ GeV-1 at 95% confidence level [2017NatPh..13..584A].
As proposed in ref. [2020JCAP...11..059G], axion helioscopes can be used to detect supernova axions in principle. One of differences between Solar axions and supernova axions is their energy. Since supernova axions are as energetic as 10–100 MeV, they are converted into $\gamma$-rays by the magnetic field. In order to detect supernova axions with axion helioscopes, one should install an additional $\gamma$-ray detector. Another difference is that, whereas the Sun is a static object, supernova explosions are transient events. It is thus necessary that the axion helioscope is pointed to the supernova event before its neutrino burst is observed. This requires the identification of the progenitor star before the explosion, which could be possible through a presupernova neutrino alert [2004APh....21..303O, 2017ApJ...851....6P, 2020ARNPS..70..121K, 2020ApJ...899..153M, 2020JCAP...05..049L, 2021NJPh...23c1201A]. It is estimated that presupernova neutrinos can provide the alert $\mathcal{O}(1)$ hours before the core bounce for a massive star closer than 1 kpc, with detectors including the Jiangmen Underground Neutrino Observatory (JUNO) [2015arXiv150807166A, 2020JCAP...05..049L], Super-Kamiokande [2003NIMPA.501..418F, 2022ApJ...935...40M, 2024ApJ...973..140A], Hyper-Kamiokande [2018arXiv180504163H], KamLAND [2014EPJC...74.3094S, 2016ApJ...818...91A, 2024ApJ...973..140A], the Deep Underground Neutrino Experiment (DUNE) [2020JInst..15.8008A], and SNO+ [2021arXiv210411687S].
The axion-photon conversion probability in an axion helioscope can estimated as [1983PhRvL..51.1415S, 1988PhRvD..37.1237R]
@@
Figure 5: The asymptotic event number of axions as a function of the axion luminosity $L_{a}$ at $t_{\mathrm{pb}}=1$ s. The left panel is based on the self-consistent simulations and the right panel is based on the post-processing. Each curve represents the results with different distances to the supernova event. The IAXO upgraded scenario is assumed as the detector. The vertical line at $L_{a}=3\times 10^{52}$ erg s-1 is an upper limit on the axion luminosity based on the SN 1987A neutrino burst [1990PhR...198....1R].
@@
where $A$ is the cross section of the helioscope, $d$ is the distance to the supernova event, $\dot{N}_{a}=L_{a}/\langle\omega\rangle$ is the axion number produced in the supernova per unit time, and $\langle\omega\rangle\approx 2.3T$ is the axion mean energy. The cross section $A$, the magnetic field $B$, and the length $L$ depend on the detectors. In this study, we adopt values for the IAXO upgraded scenario shown in ref. [2019JCAP...06..047A], where $B\sim 35$ kG, $L=22$ m, and $A=3.9$ m2.
Figure 4 shows the cumulative event number of axions from Betelgeuse, which is located at $d\approx 168$ pc [2020ApJ...902...63J]. The solid curves are the results for the models with the axion cooling effect, and the broken curves are the results obtained by the post-processing of the reference model without the axion cooling. One can find that, when the axion cooling is not considered, the event number exceeds one if $m_{a}\gtrsim 9$ meV. On the other hand, when the axion cooling is taken into account, the cumulative event number saturates earlier and its asymptotic value is significantly reduced. As a result, the expected event number does not exceed one even if axions are as heavy as $m_{a}=11$ meV. The reduction in the event number stems from the reduction in the axion luminosity, which is attributed to the lower temperature in the supernova core shown in figure 3.
@@
HD 149757 $\zeta$ Ophiuchi O 112(2) [2007AA...474..653V] 20.0 [2001MNRAS.327..353H]
HD 129056 $\alpha$ Lupi B 143(3) [2007AA...474..653V] $10.1\pm 1.0$ [2011MNRAS.410..190T]
@@
Table 2: The list of nearby stars heavier than $10M_{\odot}$ and closer than 200 pc [2020ApJ...899..153M].
@@
The axion luminosity has an upper limit based on the neutrino burst from SN 1987A. If the axion luminosity dominates over the neutrino luminosity, the duration of the SN 1987A neutrino burst should be shorter than observed. This energy-loss argument has been applied to various exotic particle models to obtain constraints on the interaction between the new particle and a Standard Model particle. In many cases, the criterion of $L_{a}<3\times 10^{52}$ erg s-1 at $t_{\mathrm{pb}}=1$ s is adopted for this purpose [e.g. 1990PhR...198....1R, 2011JCAP...01..015G, 2020JCAP...12..008L]. The vertical line in figure 5 indicates the SN 1987A limit. In our case, the models with $m_{a}\leq 9$ meV satisfies the condition, regardless of whether the axion cooling is considered or not. If we respect this upper limit on the axion luminosity, in the post-process framework, the axion signal is detectable if the star is closer than $d\sim 150$ pc. On the other hand, if we consider the axion cooling, the star should be closer than $d\sim 100$ pc for us to detect axions from it.